# Construction of 2-D Z-Complementary Array Code Sets with Flexible Lengths for Different System Requirements

### v3

@@ II-D Peak-to-Mean Envelope Power Ratio
II-E mMIMO-based Omnidirectional Beamforming
In this subsection, we introduce the concept of precoding of mMIMO with URA for omnidirectional signal transmission. As described in [13, 28], we consider a downlink transmission from a base station (BS) equipped with URA to single-antenna user equipment (UE). We consider a $P\times Q$ URA of $PQ$ antennas. Let, $A(\phi,\theta)$ be the steering matrix corresponding to the URA at the direction angle $(\phi,\theta)$, where $\phi\in[0,\pi/2]$ and $\theta\in[0,2\pi]$. Then $A(\phi,\theta)$ is given by
@@
where $\iota=\sqrt{-1}$, $10\leq i\leq P-$ 1, $10\leq j\leq Q-$ 1, $\lambda$ is the wavelength of the carrier, $d_{x}$ and $d_{y}$ are the inter-element distance of URA along the vertical and horizontal directions, respectively. At first, the data bits are modulated into a $K\times N$-STBC $\mathbf{S}$, the $(n,t)$-th entry given by $s_{n}(t)$, where $n=0,1,\dots,K-1$ and $t=0,1,\dots,N-1$ [29]. It is then beamformed with the precoding matrices $\mathbf{W}_{n}$ of size $P\times Q$ and sent over the wireless channel. The received signal in a line-of-sight (LOS) without multi-paths is given by
@@
for $10\leq t\leq N-$ 1, where $vec(\cdot)$ denotes the vector produced by vertically stacking the columns of the corresponding matrix, $(\cdot)^{T}$ denotes the transpose of the matrix ()($\cdot$)( ⋅ ), and $w(t)$ is the additive white Gaussian noise (AWGN).
@@
III Construction of 2-D ZCACS
In this section, we propose the construction of 2-D ZCACS along with its building block which is 2-D ZCAC. However, for this purpose, we first state a lemma that produces IGC code sets of flexible lengths of the form $p_{1}^{m_{1}-1}p_{2}^{m_{2}-1}\dots p_{k}^{m_{k}-1}(m_{\alpha}\geq 1,\forall\alpha)$.
III-A IGC Code Set
@@
be defined by $f=f_{1}+f_{2}+\dots+f_{k}$. Now, for $\bm{\gamma}=(\gamma_{1},\gamma_{2},\dots,\gamma_{k}),\mathbf{s}=(s_{1},s_{2},\dots,s_{k}),\mathbf{t}=(t_{1},t_{2},\dots,t_{k})\in\mathbb{Z}_{p_{1}}\times\mathbb{Z}_{p_{2}}\times\dots\times\mathbb{Z}_{p_{k}}$, we define the set of functions
@@
given by $a_{\mathbf{s},\mathbf{t}}^{\bm{\gamma}}=R_{\mathbf{t}}^{\bm{\gamma}}+T_{\mathbf{s}}$, where $R_{\mathbf{t}}^{\bm{\gamma}}:\mathbb{Z}_{p_{1}}^{m_{1}-1}\times\dots\times\mathbb{Z}_{p_{k}}^{m_{k}-1}\rightarrow\mathbb{Q}_{q}$ is defined by
@@
$\mathcal{C}_{\mathbf{s},\mathbf{t}}=\begin{bmatrix}\psi(a_{\mathbf{s},\mathbf{t}}^{\bm{\gamma}})\\ \end{bmatrix}_{\bm{\gamma}\in\mathbb{Z}_{p_{1}}\times\mathbb{Z}_{p_{2}}\times\dots\times\mathbb{Z}_{p_{k}}},$ (27)

### v4

@@ 2.4 Peak-to-Mean Envelope Power Ratio
2.5 mMIMO-based Omnidirectional Beamforming
In this subsection, discuss the application of 2-D codes in mMIMO with URA for omnidirectional signal transmission. As described in [Omni1, Omni2], we consider a downlink transmission from a base station (BS) equipped with URA to a single-antenna user equipment (UE). We consider a $P\times Q$ URA of $PQ$ antennas. Let, $A(\phi,\theta)$ be the steering matrix corresponding to the URA at the direction angle $(\phi,\theta)$, where $\phi\in[0,\pi/2]$ and $\theta\in[0,2\pi]$. Then $A(\phi,\theta)$ is given by
@@
where $\iota=\sqrt{-1}$, $0\leq i\leq P-1$, $0\leq j\leq Q-1$, $\lambda$ is the wavelength of the carrier, $d_{x}$ and $d_{y}$ are the inter-element distance of URA along the vertical and horizontal directions, respectively. At first, the data bits are modulated into a $K\times N$-STBC given by $\mathbf{S}$, where the $(n,t)$-th entry is $s_{n}(t)$, for $n=0,1,\dots,K-1$ and $t=0,1,\dots,N-1$ [STBC]. It is then beamformed with the precoding matrices $\mathbf{W}_{n}$ of size $P\times Q$ and sent over the wireless channel. The received signal in a line-of-sight (LOS) without multi-paths is given by
@@
for $0\leq t\leq N-1$, where $vec(\cdot)$ denotes the vector produced by vertically stacking the columns of the corresponding matrix, $(\cdot)^{T}$ denotes the transpose of the matrix $(\cdot)$, and $w(t)$ is the additive white Gaussian noise (AWGN).
@@
3 Construction of 2-D ZCACS
In this section, we propose the construction of 2-D ZCACS along with its constituent 2-D codes which is 2-D ZCAC. However, for this purpose, we first state a lemma that produces IGC code sets of flexible lengths of the form $p_{1}^{m_{1}-1}p_{2}^{m_{2}-1}\dots p_{k}^{m_{k}-1}(m_{\alpha}\geq 1,\forall\alpha)$.
3.1 IGC Code Set
@@
be defined by
$f=f_{1}+f_{2}+\dots+f_{k}.$ (25)
Now, for $\bm{\gamma}=(\gamma_{1},\gamma_{2},\dots,\gamma_{k}),\mathbf{s}=(s_{1},s_{2},\dots,s_{k}),\mathbf{t}=(t_{1},t_{2},\dots,t_{k})\in\mathbb{Z}_{p_{1}}\times\mathbb{Z}_{p_{2}}\times\dots\times\mathbb{Z}_{p_{k}}$, we define the set of functions
@@
given by
$a_{\mathbf{s},\mathbf{t}}^{\bm{\gamma}}=R_{\mathbf{t}}^{\bm{\gamma}}+T_{\mathbf{s}},$ (27)
where $R_{\mathbf{t}}^{\bm{\gamma}}:\mathbb{Z}_{p_{1}}^{m_{1}-1}\times\dots\times\mathbb{Z}_{p_{k}}^{m_{k}-1}\rightarrow\mathbb{Q}_{q}$ is defined by
@@
$\mathcal{C}_{\mathbf{s},\mathbf{t}}=\begin{bmatrix}\vdots\\ \psi(a_{\mathbf{s},\mathbf{t}}^{\bm{\gamma}})\\ \vdots\\ \end{bmatrix}_{\bm{\gamma}\in\mathbb{Z}_{p_{1}}\times\mathbb{Z}_{p_{2}}\times\dots\times\mathbb{Z}_{p_{k}}},$ (30)